# Learning to Learn Transferable Generative Attack for Person Re-Identification

### v1

@@ IV-C Ablation Studies
To further verify the effects of meta-learning and eliminate the effects of data zoo and model zoo, we compare with SOTA classification ensemble attacks (i.e. CWA [76], AdaEA [77], NTKL [78]). Since they only integrate multiple models without using multiple datasets, we retrained a model without the data zoo for fair comparison. As their adversarial instance perturbations cannot migrate to unseen query data, we compare the training data classification accuracy (Acc) in the cross-model setting. The results of them using the same model zoo in Tab. X show our method’s superiority and meta-learning’s effectiveness.
@@ IV-D Adversarial Example Quality
To evaluate the image quality for generated adversarial examples, we compare the SSIM [79] with other attack methods for re-id. SSIM calculates structural similarity between synthetic and natural images and larger SSIM scores indicate better quality of synthetic images. The results of SSIM between AEs(=8) and benign images on DukeMTMC are show in Tab. XI, which shows that our MTGA can obtain AEs with comparable quality.
@@ IV-E Visualization
We visualize the perturbations and adversarial examples that our MTGA generates. As Fig. 4 shows, the perturbations on adversarial examples are imperceptible. It’s hard for humans to detect the maliciously attacked adversarial examples generated by our MTGA. What’s more, the generated perturbations obtain the human shape of benign images, which indicates that our MTGA is able to understand the target that needs to be attacked and attempts to perform a full range of feature destruction for different person images, thus generating more generic adversarial attacks.
V Conclusion

### v2

@@ IV-C Ablation Studies
Method | PASS | PGFA | CLIP-ReID | |||
mAP | Rank-1 | mAP | Rank-1 | mAP | Rank-1 | |
None | 92.2 | 96.3 | 37.3 | 51.4 | 89.6 | 95.5 |
MTGA | 15.9 | 16.4 | 5.4 | 6.5 | 46.8 | 53.7 |
To further verify the effects of meta-learning and eliminate the effects of data zoo and model zoo, we compare with SOTA classification ensemble attacks (i.e., CWA [92], AdaEA [93], NTKL [94]). Since they only integrate multiple models without using multiple datasets, we retrained a model without the data zoo for fair comparison. As their adversarial instance perturbations cannot migrate to unseen query data, we compare the training data classification accuracy (Acc) in the cross-model setting. The results of them using the same model zoo in Tab. X show our method’s superiority and meta-learning’s effectiveness.
Transferability to diverse types of re-id models. To assess MTGA’s transferability to model types beyond those in the model zoo, we conducted experiments on self-supervised PASS (Market) [95] model, auxiliary-feature-enhanced PGFA (Occ-Duke) [96] model and CLIP-based CLIP-ReID (Market) [97] model. As shown in Tab. XI, MTGA significantly reduces the performance of all models, which demonstrates the MTGA’s effectiveness against diverse model types.
Analysis of key parameters. We conducted experiments on different perturbation strength, task number, learning rate, mix coefficient, PRE probability and mask percentage values. Fig. 4 shows the mDR under different settings. Larger values for perturbation strength and task number generally improve transferability, but we chose 8/255 and 5, respectively, to balance imperceptible perturbations and GPU memory. For mix coefficient and learning rate, MTGA demonstrates strong generalization and stability across different values. For PRE probability and mask percentage, the value 0 indicates that no PRE policy is performed and the attack performance is modest. When it is not 0, the effect is improved, and we choose the best parameters 0.8 and 0.2 of and as the experimental parameters.
@@ IV-D Adversarial Example Quality
To evaluate the image quality for generated adversarial examples, we compare the SSIM [98] with other attack methods for re-id. SSIM calculates structural similarity between synthetic and natural images and larger SSIM scores indicate better quality of synthetic images. The results of SSIM between AEs(=8/255) and benign images on DukeMTMC are show in Tab. XII, which shows that our MTGA can obtain AEs with comparable quality.
@@ IV-E Visualization
We visualize the perturbations and adversarial examples generated by our MTGA across multiple datasets, including Market [86], DukeMTMC [71], MSMT17 [73] and VIPeR [87]. As Fig. 5 shows, the perturbations on adversarial examples are imperceptible. It’s hard for humans to detect the maliciously attacked adversarial examples generated by our MTGA. What’s more, the generated perturbations obtain the human shape of benign images, which indicates that our MTGA is able to understand the target that needs to be attacked and attempts to perform a full range of feature destruction for different style person images, thus generating more generic adversarial attacks.
We also provide visualization of cross-model&dataset attack results by showcasing the Rank-10 matches from the target re-id models (i.e., BOT (Market) [75] and TransReID (Market) [12]) before and after applying our proposed adversarial attack on Market dataset [86], as illustrated in Fig. 6 and Fig. 7. In these figures, green boxes denote correctly matched images, red boxes indicate mismatched images, and the first column represents the query images. These visualizations demonstrate the effectiveness of our method in attacking various re-id models.
Method | Adv.Res | Randomization | JPEG | aAP | mDR |
None | 69.6 | 84.6 | 83.8 | 80.0 | - |
MetaAttack | 67.1 | 67.8 | 57.9 | 64.3 | 19.7 |
Mis-Ranking | 56.1 | 43.3 | 51.2 | 50.2 | 37.3 |
MUAP | 53.6 | 48.5 | 57.4 | 53.2 | 33.5 |
MTGA(Ours) | 40.3 | 26.3 | 31.8 | 32.8 | 59.0 |
Methods | Parameters | FLOPs | Training Time | aAP | mDR |
Baseline | |||||
MTGA |
IV-F Attack Effectiveness against Defense Method
In Tab. XIII, we present additional evaluations to assess the effectiveness of our method against various defense strategies, including adversarially trained models (Adv. Res [27]), input preprocessing techniques (JPEG compression [99]), and denoising-based methods (Randomization [100]). For the JPEG defense, a compression rate of 60% was applied, and the victim model used for evaluation was BOT (Market) [75]. Our method consistently demonstrates superior attack effectiveness across these defenses. Notably, against these three categories of defenses, our approach achieves an mDR of 59.0%, highlighting the pressing need for more robust defense mechanisms to ensure the security of re-ID systems.
V Discussion
Computation cost. The comparison of computational cost and transferability performance between the baseline model and our proposed MTGA is presented in Tab. XIV. For attack inference, both methods utilize the same adversarial generator, resulting in identical model parameters and FLOPs. Regarding training time, our MTGA incorporates a meta-learning scheme, which increases the training time to more than twice that of the baseline without meta-learning optimization. However, this additional training time yields a significant improvement in transferability performance. Importantly, during the testing phase, both methods maintain identical inference efficiency, ensuring no additional computational overhead.
Impact. Our proposed adversarial attack method may potentially be exploited by malicious attackers to compromise surveillance systems, which also alters the security of re-id system and provides an effective benchmark for testing the robustness of real-world models. In the future, we plan to leverage the adversarial examples proposed in this work to further explore the development of more robust re-id models.
Limitation and future work. Our method integrates data zoo and model zoo with meta-learning, which is time-consuming and resource-intensive in training process. To address this limitation, future work will explore the introduction of visual-language models, which can provide joint visual and text representations, enabling more efficient and effective feature disruption while reducing computational costs.
VI Conclusion